Find the molar mass of an unknown gas which diffuses twice as fast as Cl2 gas.


1
Expert's answer
2020-12-01T09:19:26-0500


Q145868

Find the molar mass of an unknown gas which diffuses twice as fast as Cl2 gas.


Solution:


Graham’s Law: Rate of diffusion or of effusion of a gas is inversely proportional to the square root of its molar mass.


For two gases A and B, their rate of diffusion and molar mass can be related by the formula


"{{Rate_A}\over{Rate _B}} = \sqrt({ {M_B}\over{M_A} })"



where, Rate A = rate of diffusion of gas A.

MA = molar mass of gas A.

Rate B = rate of diffusion of gas B

MB = molar mass of gas B.


Suppose the unknown gas is ‘X’ and the molar mass of the unknown gas is MX.


Molar mass of Cl2 = 2 * atomic mass of Cl = 2 * 35.453 g/mol

= 70.906 g/mol


In Question, we are given that the diffusion rate of the unknown gas is twice as fast as Cl2 gas.

Which means

Rate of diffusion of unknown gas = 2 * Rate of diffusion of Cl2 .


Rate X = 2 * Rate Cl2 .


which can also be written as


"{{Rate_X}\over{Rate _{Cl2}}} = 2;"


plug this and the molar mass of Cl2 in Graham’s equation, we have  


"{{Rate_X}\over{Rate _{Cl2}}} = \sqrt({ {M_{Cl2}}\over{M_X} }) ;"


"2 = \sqrt({ {70.906g/mol}\over{M_X} }) ;"


squaring both sides of the equation, we have  


"4 ={ {70.906g/mol}\over{M_X} } ;"


which can also be written as  


"M_X ={ {70.906g/mol}\over{4} } ;"


"M_X = 17.73g\/mol"


Hence the molar mass of the unknown gas is 17.73g/mol.